OpenStudy (anonymous):

Line IJ has an equation of a line y = −3x − 8. Which of the following could be an equation for a line that is parallel to line IJ?

OpenStudy (anonymous):

A. y = 3x + 7 B. y = 1 over 3x + 7 C. y = −3x + 7 D. y = −1 over 3x +

OpenStudy (midhun.madhu1987):

The line which have slope same as that of line IJ will be parallel to IJ.

OpenStudy (midhun.madhu1987):

The quation of line is given by y = mx + c where m = slope

OpenStudy (anonymous):

ok

OpenStudy (midhun.madhu1987):

y = −3x − 8 what is the slope of this line?

OpenStudy (anonymous):

x=3, y=8

OpenStudy (midhun.madhu1987):

y = mx + c where m = slope y = −3x − 8 So, what is the value of m?

OpenStudy (anonymous):

3

OpenStudy (midhun.madhu1987):

Nope... it is -3 So from 4 options which equation has same slope as that of IJ?

OpenStudy (hybrik):

You can also just find a formula with the exact same slope

OpenStudy (hybrik):

[-3]x-8 [-3] is the same slope

OpenStudy (anonymous):

It C thank you